Please help me with these two questions I am highly confused.

1.A mixture of argon and helium gases is maintained in a 6.53 L flask at a pressure of 2.71 atm and a temperature of 25 °C. If the gas mixture contains 10.7 grams of argon, the number of grams of helium in the mixture is______ g.

2. The stopcock connecting a 4.10 L bulb containing argon gas at a pressure of 9.21 atm, and a 8.33 L bulb containing oxygen gas at a pressure of 2.24 atm, is opened and the gases are allowed to mix. Assuming that the temperature remains constant, the final pressure in the system is ________atm.
Answer #1

* N = 6.53L, P=2011 atin T=25+213 = 298k Ar & Hen. Non reached gases. R=0.0821 h atin key molt :- PV=NTR (2.11 alfon) (6.534)Pressure due to O. P, Q, Pave P = 2.auatur P. ₂ 28.832 Rig 2 19.432 (2.24 atm (8.334 - P. (120434) 2 = 10 501 alin . final pr